The GNU Shepherd Manual

For use with the GNU Shepherd 0.10.4 Last updated 31 March 2024

Wolfgang Jährling Ludovic Courtès

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2002, 2003 Wolfgang Jährling Copyright \begin{matrix}c \\ ◯ \end{matrix} 2013, 2016, 2018, 2019, 2020, 2022, 2023 Ludovic Court^^c3^^a8s Copyright \begin{matrix}c \\ ◯ \end{matrix} 2020 Brice Waegeneire Copyright \begin{matrix}c \\ ◯ \end{matrix} 2020 Oleg Pykhalov Copyright \begin{matrix}c \\ ◯ \end{matrix} 2020, 2023 Jan (janneke) Nieuwenhuizen

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Jump Start. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 herd and shepherd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.1 Invoking shepherd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Invoking herd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3 Invoking reboot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.4 Invoking halt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4.1 Defining Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2 Service Registry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.3 Interacting with Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.4 Service De- and Constructors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.5 The root Service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.6 Legacy GOOPS Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 4.7 Service Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.8 Managing User Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

5 Service Collection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

5.1 Monitoring Service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5.2 Read-Eval-Print Loop Service. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

6 Misc Facilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

6.1 Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 6.2 Communication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

7 Internals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

7.1 Coding Standards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7.2 Design Decisions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7.3 Service Internals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

Appendix A GNU Free Documentation License . . 30

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

Procedure and Macro Index . . . . . . . . . . . . . . . . . . . . . . 40 Variable Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

Type Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

1 Introduction

This manual documents the GNU Daemon Shepherd, or GNU Shepherd for short. The Shepherd looks after system services, typically daemons. It is used to start and stop them in a reliable fashion. For instance, it will dynamically determine and start any other services that our desired service depends upon.

The Shepherd is the init system of the GNU operating system—it is the first user process that gets started, typically with PID 1, and runs as root. Normally the purpose of init systems is to manage all system-wide services, but the Shepherd can also be a useful tool assisting unprivileged users in the management of their own daemons.

Flexible software requires some time to master and the Shepherd is no different. But don’t worry: this manual should allow you to get started quickly. Its first chapter is designed as a practical introduction to the Shepherd and should be all you need for everyday use (see Chapter 2 [Jump Start], page 2). In chapter two we will describe the herd and shepherd programs, and their relationship, in more detail (Chapter 3 [herd and shepherd], page 5). Subsequent chapters provide a full reference manual and plenty of examples, covering all of Shepherd’s capabilities. Finally, the last chapter provides information for those souls brave enough to hack the Shepherd itself.

The Shepherd was formerly known as “dmd”, which stands for Daemon Managing Daemons (or Daemons-Managing Daemon?).

This program is written in Guile Scheme. Guile is also the Shepherd’s configuration language. See Section “Introduction” in GNU Guile Reference Manual, for an introduction to Guile. We have tried to make the Shepherd’s basic features as accessible as possible—you should be able to use these even if you do not know how to program in Scheme. A basic grasp of Guile is required only if you wish to make use of the Shepherd’s more advanced features.

2 Jump Start

This chapter gives a short overview of the Shepherd. It is enough if you just need the basic features of it. As it is not assumed that readers are familiar with all the involved issues, a very experienced user might be annoyed by the often very detailed descriptions in this introduction. Those users are encouraged to just skip to the reference section.

Note that all the full file names in the following text are based on the assumption that you have installed the Shepherd with an empty prefix. If your Shepherd installation for example resides in /usr/local instead, add this directory name in front of the absolute file names mentioned below.

When shepherd gets started, it reads and evaluates a configuration file. When it is started with superuser privileges, it tries to use /etc/shepherd.scm. When started as

normal user, it looks for a file called $XDG_CONFIG_HOME/shepherd/init.scm. If the XDG_
CONFIG_HOME environment variable is not defined, $HOME/.config/shepherd/init.scm is

used instead (see ⟨undefined⟩ [Managing User Services ], page ⟨undefined⟩). With the option --config (or, for short, -c), you can specify where to look instead. So if you want to start shepherd with an alternative file, use one of the following commands:

     shepherd --config=/etc/shepherd.scm.old
     shepherd -c /etc/shepherd.scm.old

As the final “d” suggests, shepherd is just a daemon that (usually) runs in the background, so you will not interact with it directly. After it is started, shepherd will listen on a socket special file, usually /var/run/shepherd/socket, for further commands. You use the tool herd to send these commands to shepherd. Usage of herd is simple and straightforward: To start a service called apache, you use:

     herd start apache

When you do this, all its dependencies will get resolved. For example, a webserver is quite likely to depend on working networking, thus it will depend on a service called networking. So if you want to start apache, and networking is not yet running, it will automatically be started as well. The current status of all the services defined in the configuration file can be queried like this:

     herd status

Or, to get additional details about each service, run:

     herd detailed-status

In this example, this would show the networking and apache services as started. If you just want to know the status of the apache service, run:

     herd status apache

You may also view a log of service events, including the time at which each service was started or stopped, by running:

     herd log

Services and their dependencies form a graph, which you can view, for instance with the

help of xdot (https://github.com/jrfonseca/xdot.py), by running:
     herd graph | xdot -

You can stop a service and all the services that depend on it will be stopped. Using the example above, if you stop networking, the service apache will be stopped as well—which makes perfect sense, as it cannot work without the network being up. To actually stop a service, you use the following, probably not very surprising, command:

     herd stop networking

There are two more actions you can perform on every service: the actions enable and disable are used to prevent and allow starting of the particular service. If a service is intended to be restarted whenever it terminates (how this can be done will not be covered in this introduction), but it is respawning too often in a short period of time (by default 5 times in 5 seconds), it will automatically be disabled. After you have fixed the problem that caused it from being respawned too fast, you can start it again with the commands:

     herd enable foo
     herd start foo

But there is far more you can do than just that. Services can not only simply depend on other services, they can also depend on virtual services. A virtual service is a service that is provided by one or more services additionally. For instance, a service called exim might provide the virtual service mailer-daemon. That could as well be provided by a service called smail, as both are mailer-daemons. If a service needs any mailer-daemon, no matter which one, it can just depend on mailer-daemon, and one of those who provide it gets started (if none is running yet) when resolving dependencies. The nice thing is that, if trying to start one of them fails, shepherd will go on and try to start the next one, so you can also use virtual services for specifying fallbacks.

Additionally to all that, you can perform service-specific actions. Coming back to our original example, apache is able to reload its modules, therefore the action reload-modules might be available:

     herd reload-modules apache

Service-specific actions can only be used when the service is started, i.e. the only thing you can do to a stopped service is starting it. An exception exists, see below.

There are two actions which are special, because even if services can implement them on their own, a default implementation is provided by shepherd (another reason why they are special is that the default implementations can be called even when the service is not running; this inconsistency is just to make it more intuitive to get information about the status of a service, see below).

These actions are restart and status. The default implementation of restart calls stop and start on the affected service, taking care to also restart any dependent services. The default implementation of status displays some general information about the service, like what it provides, and what it depends on.

A special service is root, which is used for controlling the Shepherd itself. You can also reference to this service as shepherd. It implements various actions. For example, the status action displays which services are started and which ones are stopped, whereas detailed-status has the effect of applying the default implementation of status to all services one after another. The load action is unusual insofar as it shows a feature that is actually available to all services, but which we have not seen yet: It takes an additional argument. You can use load to load arbitrary code into the Shepherd at runtime, like this:

     herd load shepherd ~/additional-services.scm

In the same vein the special action doc describes its service when called without an argument or describes a service-specific action when called with the action as the additional arguments. You can even get the list of the service-specific actions a service provides when using with the additional argument list-actions.

$ herd doc root
The root service is used to operate on shepherd itself.
$ herd doc root list-actions
root (help status halt power-off load eval unload reload daemonize cd restart)
$ herd doc root action power-off
power-off: Halt the system and turn it off.

This is enough now about the herd and shepherd programs, we will now take a look at how to configure the Shepherd. In the configuration file, we need mainly the definition of services. We can also do various other things there, like starting a few services already.

FIXME: Finish. For now, look at the doc/examples/ subdirectory.

...

Ok, to summarize:

• shepherd is a daemon, herd the program that controls it.

• You can start, stop, restart, enable and disable every service, as well as display its status.

• You can perform additional service-specific actions, which you can also list.

• Actions can have arguments.

• You can display the status of a service, even if the service does not provide a specific implementation for this action. The same is true for restarting.

• The root/shepherd service is used to control shepherd itself.

3 herd and shepherd

The daemon that runs in the background and is responsible for controlling the services is shepherd, while the user interface tool is called herd: it’s the command that allows you to actually herd your daemons^{1}. To perform an action, like stopping a service or calling an action of a service, you use the herd program. It will communicate with shepherd over a Unix Domain Socket.

Thus, you start shepherd once, and then always use herd whenever you want to do something service-related. Since herd passes its current working directory to shepherd, you can pass relative file names without trouble. Both shepherd and herd understand the

standard arguments --help, --version and --usage.

3.1 Invoking shepherd

The shepherd program has the following synopsis:

shepherd [option...]

It accepts the following options:

‘-c file’
‘--config=file’

Read and evaluate file as the configuration script on startup. Scheme code in file is evaluated in the context of a fresh module where bindings from the (shepherd service) module are available, in addition to the default set of Guile bindings. It may typically perform three actions:

1. defining services using the service procedure (see Section 4.1 [Defining Services], page 8);

2. registering those services with register-services (see Section 4.2 [Service Registry], page 12);

3. starting some or all of those services with start-in-the-background (see Section 4.3 [Interacting with Services], page 12).

See Section 4.7 [Service Examples], page 20, for examples of what file might look like. Note that file is evaluated asynchronously: shepherd may start listening for client connections (with the herd command) before file has been fully loaded. Errors in file such as service startup failures or uncaught exceptions do not cause shepherd to stop. Instead the error is reported, leaving users the opportunity to inspect service state and, for example, to load an updated config file with:

     herd load root file
‘-I’
‘--insecure’

Do not check if the directory where the socket—our communication rendezvous with herd—is located has permissions 700. If this option is not specified, shepherd will abort if the permissions are not as expected.

1 In the past, when the GNU Shepherd was known as GNU dmd, the herd command was called deco, for

DaEmon COntroller.

‘-l [file]’
‘--logfile[=file]’

Log output into file. For unprivileged users, the default log file is $XDG_STATE_

HOME/shepherd/shepherd.log with $XDG_STATE_HOME defaulting to
$HOME/.local/state.

When running as root, the default behavior is to connect to /dev/log, the syslog socket (see Section “Overview of Syslog” in The GNU C Library Reference Manual). A syslog daemon, syslogd, is expected to read messages from there (see Section “syslogd invocation” in GNU Inetutils). When /dev/log is unavailable, for instance because syslogd is not running, as is the case during system startup and shutdown, shepherd falls back to the Linux kernel ring buffer, /dev/kmsg. If /dev/kmsg is missing, as is the case on other operating systems, it falls back to /dev/console.

‘--pid[=file]’

When shepherd is ready to accept connections, write its PID to file or to the standard output if file is omitted.

‘-s file’
‘--socket=file’

Receive further commands on the socket special file file. If this option

is not specified, localstatedir/run/shepherd/socket is taken when

running as root; when running as an unprivileged user, shepherd listens to /run/user/uid/shepherd/socket, where uid is the user’s numerical ID^{2}, or

to $XDG_RUNTIME_DIR/shepherd when the XDG_RUNTIME_DIR environment

variable is defined. If - is specified as file name, commands will be read from standard input, one per line, as would be passed on a herd command line (see Section 3.2 [Invoking herd], page 6).

‘--quiet’ Synonym for --silent.

3.2 Invoking herd

The herd command is a generic client program to control a running instance of shepherd (see Section 3.1 [Invoking shepherd], page 5). When running as root, it communicates with the system instance—the process with PID 1; when running as a normal user, it communicates with the user’s instance, which is a regular, unprivileged process managing the user’s own services. For example, the following command displays the status of all the system services:

sudo herd status

Conversely, the command below displays the status of user services, assuming a user shepherd is running:

herd status

2 On GNU/Linux, the /run/user/uid directory is typically created by elogind or by systemd, which are

available in most distributions.

The command has the following synopsis:

     herd [option...] action [service [arg...]]

It causes the action of the service to be invoked. When service is omitted and action is status or detailed-status, the root service is used^{3} (see Section 4.5 [The root Service], page 18, for more information on the root service.)

For each action, you should pass the appropriate args. Actions that are available for

every service are start, stop, restart, status, enable, disable, and doc.

If you pass a file name as an arg, it will be passed as-is to the Shepherd, thus if it is not an absolute name, it is local to the current working directory of shepherd, not to herd.

The herd command understands the following option:

‘-s file’
‘--socket=file’

Send commands to the socket special file file. If this option is not specified,

          localstatedir/run/shepherd/socket is taken.

The herd command returns zero on success, and a non-zero exit code on failure. In particular, it returns a non-zero exit code when action or service does not exist and when the given action failed.

3.3 Invoking reboot

The reboot command is a convenience client program to instruct the Shepherd (when used as an init system) to stop all running services and reboot the system. It has the following synopsis:

     reboot [option...]

It is equivalent to running herd stop shepherd. The reboot command understands the following option:

‘-s file’
‘--socket=file’

Send commands to the socket special file file. If this option is not specified,

          localstatedir/run/shepherd/socket is taken.

3.4 Invoking halt

The halt command is a convenience client program to instruct the Shepherd (when used as an init system) to stop all running services and turn off the system. It has the following synopsis:

     halt [option...]

It is equivalent to running herd power-off shepherd. As usual, the halt command understands the following option:

‘-s file’
‘--socket=file’

Send commands to the socket special file file. If this option is not specified,

          localstatedir/run/shepherd/socket is taken.

3 This shorthand does not work for other actions such as stop, because inadvertently typing herd stop

would stop all the services, which could be pretty annoying.

4 Services

The service is obviously a very important concept of the Shepherd. On the Guile level, a service is represented as a record of type <service>. Each service has a number of properties that you specify when you create it: how to start it, how to stop it, whether to respawn it when it stops prematurely, and so on. At run time, each service has associated state: whether it is running or stopped, what PID or other value is associated with it, and so on.

This section explains how to define services and how to query their configuration and state using the Shepherd’s programming interfaces.

Note: The programming interface defined in this section may only be used within a shepherd process. Examples where you may use it include:

• the shepherd configuration file (see Section 4.7 [Service Examples], page 20);

• as an argument to herd eval root ... (see Section 4.5 [The root Service], page 18);

• at the REPL (see Section 5.2 [REPL Service], page 23).

These procedures may not be used in Guile processes other than shepherd itself.

4.1 Defining Services

A service is created by calling the service procedure, from the (shepherd service) module (automatically visible from your configuration file), as in this example:

     (service
       '(sshd ssh-daemon)  ;for convenience, give it two names
       #:start (make-forkexec-constructor
                '("/usr/sbin/sshd" "-D")
                #:pid-file "/etc/ssh/sshd.pid")
       #:stop (make-kill-destructor)
       #:respawn? #t)

The example above creates a service with two names, sshd and ssh-daemon. It is started by invoking /usr/sbin/sshd, and it is considered up and running as soon as its PID file /etc/ssh/sshd.pid is available. It is stopped by terminating the sshd process. Finally, should sshd terminate prematurely, it is automatically respawned. We will look at #:start and #:stop later (see Section 4.4 [Service De- and Constructors], page 14), but first, here is the reference of the service procedure and its optional keyword arguments.

[Procedure] service provision [#:requirement ’()] [#:one-shot? #f] [#:transient? #f] [#:respawn? #f] [#:start (const #t)] [#:stop (const #f)] [#:actions (actions)] [#:termination-handler default-service-termination-handler] [#:documentation #f]

Return a new service with the given provision, a list of symbols denoting what the service provides. The first symbol in the list is the canonical name of the service, thus it must be unique.

The meaning of keyword arguments is as follows:

#:requirement

#:requirement is, like provision, a list of symbols that specify services. In this case, they name what this service depends on: before the service can be started, services that provide those symbols must be started. Note that every name listed in #:requirement must be registered so it can be resolved (see Section 4.2 [Service Registry], page 12).

#:respawn?

Specify whether the service should be respawned by shepherd. If this slot has the value #t, then, assuming the service has an associated process (its “running value” is a PID), restart the service if that process terminates. There is a limit to avoid endless respawning: when the service gets respawned “too fast”, it is disabled—see #:respawn-limit below.

#:respawn-delay

Specify the delay before a service is respawned, in seconds (including a fraction), for services marked with #:respawn? #t. Its default value is (default-respawn-delay) (see Section 4.4 [Service De- and Constructors], page 14).

#:respawn-limit

Specify the limit that prevents shepherd from respawning too quickly the service marked with #:respawn? #t. Its default value is (default-respawn-limit) (see Section 4.4 [Service De- and Constructors], page 14). The limit is expressed as a pair of integers: the first integer, n, specifies a number of consecutive respawns and the second integer, t, specifies a number of seconds. If the service gets respawned more than n times over a period of t seconds, it is automatically disabled (see Section 4.3 [Interacting with Services], page 12). Once it is disabled, the service must be explicitly re-enabled using herd enable service before it can be started again. Consider the service below:

     (service '(xyz)
              #:start (make-forkexec-constructor ...)
              #:stop (make-kill-destructor)
              #:respawn? #t
              #:respawn-limit '(3 . 5))

The effect is that this service will be respawned at most 3 times over a period of 5 seconds; if its associated process terminates a fourth time during that period, the service will be marked as disabled.

#:one-shot?

Whether the service is a one-shot service. A one-shot service is a service that, as soon as it has been successfully started, is marked as “stopped.” Other services can nonetheless require one-shot services. One-shot services are useful to trigger an action before other services are started, such as a cleanup or an initialization action.

As for other services, the start method of a one-shot service must return a truth value to indicate success, and false to indicate failure.

#:transient?

Whether the service is a transient service. A transient service is automatically unregistered when it terminates, be it because its stop method is called or because its associated process terminates.

This is useful in the uncommon case of synthesized services that may not be restarted once they have completed.

#:start Specify the constructor of the service, which will be called to start the service. This must be a procedure that accepts any number of arguments; those arguments will be those supplied by the user, for instance by passing them to herd start. If the starting attempt failed, it must return #f or throw an exception; otherwise, the return value is stored as the running value of the service.

Note: Constructors must terminate, successfully or not, in a timely fashion, typically less than a minute. Failing to do that, the service would remain in “starting” state and users would be unable to stop it.

See Section 4.4 [Service De- and Constructors], page 14, for info on common service constructors.

#:stop This is the service destructor: a procedure of one or more arguments that should stop the service. It is called whenever the user explicitly stops the service; its first argument is the running value of the service, subsequent arguments are user-supplied. Its return value will again be stored as the running value, so it should return #f if it is now possible again to start the service at a later point.

Note: Destructors must also terminate in a timely fashion, typically less than a minute. Failing to do that, the service would remain in “stopping” state and users would be unable to stop it.

See Section 4.4 [Service De- and Constructors], page 14, for info on common service destructors.

#:termination-handler

The procedure to call when the process associated with the service terminates. It is passed the service, the PID of the terminating process, and its exit status, an integer as returned by waitpid (see Section “Processes” in GNU Guile Reference Manual).

The default handler is the default-service-termination-handler pro-

cedure, which respawns the service if applicable.

#:actions

The additional actions that can be performed on the service when it is running. A typical example for this is the restart action. The actions macro can be used to defined actions (see below).

A special service that every other service implicitly depends on is the root (also known as shepherd) service. See Section 4.5 [The root Service], page 18, for more information.

Services and their dependencies form a graph. At the command-line, you can view that export a representation of that graph that can be consumed by any application that understands the Graphviz format, such as xdot (https://github.com/jrfonseca/xdot.

py):
herd graph | xdot -

Service actions are defined using the actions macro, as shown below.

[Macro] actions (name proc) …

Create a value for the #:actions parameter of service. Each name is a symbol and each proc the corresponding procedure that will be called to perform the action. A proc has one argument, which will be the running value of the service.

Naturally, the (shepherd service) provides procedures to access this information for a given service object:

[Procedure] service-provision service

Return the symbols provided by service.

[Procedure] service-canonical-name service

Return the canonical name of service, which is the first element of the list returned

by service-provision.
[Procedure] service-requirement service

Return the list of services required by service as a list of symbols.

[Procedure] one-shot-service? service
[Procedure] transient-service? service

Return true if service is a one-shot/transient service.

[Procedure] respawn-service? service

Return true if service is meant to be respawned if its associated process terminates prematurely.

[Procedure] service-respawn-delay service

Return the respawn delay of service, in seconds (an integer or a fraction or inexact number). See #:respawn-delay above.

[Procedure] service-respawn-limit service

Return the respawn limit of service, expressed as a pair—see #:respawn-limit above.

[Procedure] service-documentation service

Return the documentation (a string) of service.

4.2 Service Registry

At run time, shepherd maintains a service registry that maps service names to service records. Service dependencies are expressed as a list of names passed as #:requirement to the service procedure (see Section 4.1 [Defining Services], page 8); these names are looked up in the registry. Likewise, when running herd start sshd or similar commands (see Chapter 2 [Jump Start], page 2), the service name is looked up in the registry.

Consequently, every service must be appear in the registry before it can be used. A typical configuration file thus includes a call to the register-services procedure (see Section 4.7 [Service Examples], page 20). The following procedures let you interact with the registry.

[Procedure] register-services services

Register services so that they can be looked up by name, for instance when resolving dependencies. Each name uniquely identifies one service. If a service with a given name has already been registered, arrange to have it replaced when it is next stopped. If it is currently stopped, replace it immediately.

[Procedure] unregister-services services

Remove all of services from the registry, stopping them if they are not already stopped.

[Procedure] lookup-service name

Return the service that provides name, #f if there is none.

[Procedure] for-each-service proc

Call proc, a procedure taking one argument, once for each registered service.

[Procedure] lookup-running name

Return the running service that provides name, or false if none.

4.3 Interacting with Services

What we have seen so far is the interface to define a service and to access it (see Section 4.1 [Defining Services], page 8). The procedures below, also exported by the (shepherd service) module, let you modify and access the state of a service. You may use them in your configuration file, for instance to start some or all of the services you defined (see Section 4.7 [Service Examples], page 20).

Under the hood, each service record has an associated fiber (really: an actor) that encapsulates its state and serves user requests—a fiber is a lightweight execution thread (see Section 7.3 [Service Internals], page 29).

The procedures below let you change the state of a service.

[Procedure] start-service service . args

Start service and its dependencies, passing args to its start method. Return its running value, #f on failure.

[Procedure] stop-service service . args

Stop service and any service that depends on it. Return the list of services that have been stopped (including transitive dependent services). If service is not running, print a warning and return its canonical name in a list.

[Procedure] perform-service-action service the-action . args

Perform the-action (a symbol such as 'restart or 'status) on service, passing it args. The meaning of args depends on the action.

The start-in-the-background procedure, described below, is provided for your convenience: it makes it easy to start a set of services right from your configuration file, while letting shepherd run in the background.

[Procedure] start-in-the-background services

Start the services named by services, a list of symbols, in the background. In other words, this procedure returns immediately without waiting until all of services have been started.

This procedure can be useful in a configuration file because it lets you interact right away with shepherd using the herd command.

The following procedures let you query the current state of a service.

[Procedure] service-running? service
[Procedure] service-stopped? service
[Procedure] service-enabled? service

Return true if service is currently running/stopped/enabled, false otherwise.

[Procedure] service-status service

Return the status of service as a symbol, one of: 'stopped, 'starting, 'running,

or 'stopping.
[Procedure] service-running-value service

Return the current “running value” of service—a Scheme value associated with it. It is #f when the service is stopped; otherwise, it is a truth value, such as an integer denoting a PID (see Section 4.4 [Service De- and Constructors], page 14).

[Procedure] service-status-changes service

Return the list of symbol/timestamp pairs representing recent state changes for service.

[Procedure] service-startup-failures service
[Procedure] service-respawn-times service

Return the list of startup failure times or respawn times of service.

[Procedure] service-replacement service

Return the replacement of service, or #f if there is none.

The replacement is the service that will replace service when it is eventually stopped.

See Section 7.3 [Service Internals], page 29, if you’re curious about the nitty-gritty details!

4.4 Service De- and Constructors

Each service has a start procedure and a stop procedure, also referred to as its constructor and destructor (see Chapter 4 [Services], page 8). The procedures listed below return procedures that may be used as service constructors and destructors. They are flexible enough to cover most use cases and carefully written to complete in a timely fashion.

[Procedure] make-forkexec-constructor command [#:user #f] [#:group #f]

[#:supplementary-groups ’()] [#:pid-file #f] [#:pid-file-timeout (default-pid-file-timeout)] [#:log-file #f] [#:directory (default-service-directory)] [#:file-creation-mask #f] [#:create-session? #t] [#:resource-limits ’()] [#:environment-variables (default-environment-variables)]

Return a procedure that forks a child process, closes all file descriptors except the standard output and standard error descriptors, sets the current directory to directory, sets the umask to file-creation-mask unless it is #f, changes the environment to environment-variables (using the environ procedure), sets the current user to user the current group to group unless they are #f and supplementary groups to supplementary-groups unless they are '(), and executes command (a list of strings.) When create-session? is true, the child process creates a new session with setsid and becomes its leader. The result of the procedure will be the PID of the child process.

Note: This will not work as expected if the process “daemonizes” (forks); in that case, you will need to pass #:pid-file, as explained below.

When pid-file is true, it must be the name of a PID file associated with the process being launched; the return value is the PID once that file has been created. If pid-file does not show up in less than pid-file-timeout seconds, the service is considered as failing to start.

When log-file is true, it names the file to which the service’s standard output and standard error are redirected. log-file is created if it does not exist, otherwise it is appended to.

Guile’s setrlimit procedure is applied on the entries in resource-limits. For example, a valid value would be:

     '((nproc 10 100)         ;number of processes
       (nofile 4096 4096))    ;number of open file descriptors

[Procedure] make-kill-destructor [signal] [#:grace-period

(default-process-termination-grace-period)]

Return a procedure that sends signal to the process group of the PID given as argument, where signal defaults to SIGTERM. If the process is still running after graceperiod seconds, send it SIGKILL. The procedure returns once the process has terminated.

This does work together with respawning services, because in that case the stop method of the <service> arranges so that the service is not respawned.

The make-forkexec-constructor procedure builds upon the following procedures.

[Procedure] exec-command command [#:user #f] [#:group #f] [#:supplementary-groups ’()] [#:log-file #f] [#:log-port #f] [#:input-port #f] [#:directory (default-service-directory)] [#:file-creation-mask #f] [#:create-session? #t] [#:resource-limits ’()] [#:environment-variables (default-environment-variables)]

[Procedure] fork+exec-command command [#:user #f] [#:group #f] [#:supplementary-groups ’()] [#:log-file #f] [#:log-encoding "UTF-8"] [#:directory (default-service-directory)] [#:file-creation-mask #f] [#:create-session? #t] [#:resource-limits ’()] [#:environment-variables (default-environment-variables)]

Run command as the current process from directory, with file-creation-mask if it’s true, with rlimits, and with environment-variables (a list of strings like "PATH=/bin".) File descriptors 1 and 2 are kept as is or redirected to either log-port or log-file if it’s true, whereas file descriptor 0 (standard input) points to input-port or /dev/null; all other file descriptors are closed prior to yielding control to command. When createsession? is true, call setsid first (see Section “Processes” in GNU Guile Reference Manual). By default, command is run as the current user. If the user keyword argument is present and not false, change to user immediately before invoking command. user may be a string, indicating a user name, or a number, indicating a user ID. Likewise, command will be run under the current group, unless the group keyword argument is present and not false, and supplementary-groups is not '(). fork+exec-command does the same as exec-command, but in a separate process whose PID it returns.

[Variable] default-environment-variables

This parameter (see Section “Parameters” in GNU Guile Reference Manual) specifies the default list of environment variables to be defined when the procedures above create a new process. It must be a list of strings where each string has the format name=value. It defaults to what environ returns when the program starts (see Section “Runtime Environment” in GNU Guile Reference Manual).

[Variable] default-pid-file-timeout

This parameter (see Section “Parameters” in GNU Guile Reference Manual) specifies the default PID file timeout in seconds, when #:pid-file is used (see above). It defaults to 5 seconds.

[Variable] default-process-termination-grace-period

This parameter (see Section “Parameters” in GNU Guile Reference Manual) specifies the “grace period” (in seconds) after which a process that has been sent SIGTERM or some other signal to gracefully exit is sent SIGKILL for immediate termination. It defaults to 5 seconds.

[Variable] default-respawn-delay

This parameter specifies the default value of the #:respawn-delay parameter of service (see Section 4.1 [Defining Services], page 8). It defaults to 0.1, meaning a 100ms delay before respawning a service.

[Variable] default-respawn-limit

This parameter specifies the default value of the #:respawn-limit parameter of service (see Section 4.1 [Defining Services], page 8). As an example, suppose you add this line to your configuration file:

     (default-respawn-limit '(3 . 10))

The effect is that services will be respawned at most 3 times over a period of 10 seconds before being disabled.

One may also define services meant to be started on demand. In that case, shepherd listens for incoming connections on behalf of the program that handles them; when it accepts an incoming connection, it starts the program to handle them. The main benefit is that such services do not consume resources until they are actually used, and they do not slow down startup.

These services are implemented following the protocol of the venerable inetd “super server” (see Section “inetd invocation” in GNU Inetutils). Many network daemons can be invoked in “inetd mode”; this is the case, for instance, of sshd, the secure shell server of the OpenSSH project. The Shepherd lets you define inetd-style services, specifically those in nowait mode where the daemon is passed the newly-accepted socket connection while shepherd is in charge of listening.

Listening endpoints for such services are described as records built using the endpoint procedure:

[Procedure] endpoint address [#:name "unknown"] [#:style SOCK STREAM] [#:backlog 128] [#:socket-owner (getuid)] [#:socket-group (getgid)] [#:socket-directory-permissions #o755] [#:bind-attempts (default-bind-attempts)]

Return a new endpoint called name of address, an address as return by make-socket-address, with the given style and backlog. When address is of type AF_INET6, the endpoint is IPv6-only. Thus, if you want a service available both on IPv4 and IPv6, you need two endpoints. For example, below is a list of endpoints to listen on port 4444 on all the network interfaces, both in IPv4 and IPv6 (“0.0.0.0” for IPv4 and “::0” for IPv6):

     (list (endpoint (make-socket-address AF_INET INADDR_ANY 4444))
           (endpoint (make-socket-address AF_INET6 IN6ADDR_ANY 4444)))
This is the list you would pass to make-inetd-constructor or make-systemd-
constructor—see below.

When address is of type AF_UNIX, socket-owner and socket-group are strings or integers that specify its ownership and that of its parent directory; socket-directory-permissions specifies the permissions for its parent directory. Upon ‘EADDRINUSE’ (“Address already in use”), up to bind-attempts attempts will be made to bind on address, one every second.

[Variable] default-bind-attempts

This parameter specifies the number of times, by default, that shepherd will try to bind an endpoint address if it happens to be already in use.

The inetd service constructor takes a command and a list of such endpoints:

[Procedure] make-inetd-constructor command endpoints

[#:service-name-stem ] [#:requirements ’()] [#:max-connections (default-inetd-max-connections)] [#:user #f] [#:group #f] [#:supplementary-groups ’()] [#:directory (default-service-directory)] [#:file-creation-mask #f] [#:create-session? #t] [#:resource-limits ’()] [#:environment-variables (default-environment-variables)]

Return a procedure that opens sockets listening to endpoints, a list of objects as returned by endpoint, and accepting connections in the background. Upon a client connection, a transient service running command is spawned. Only up to max-connections simultaneous connections are accepted; when that threshold is reached, new connections are immediately closed. The remaining arguments are as for make-forkexec-constructor.

[Procedure] make-inetd-destructor

Return a procedure that terminates an inetd service.

The last type is systemd-style services. Like inetd-style services, those are started on demand when an incoming connection arrives, but using the protocol devised by the systemd service manager and referred to as socket activation (https://www.freedesktop.

org/software/systemd/man/daemon.html#Socket-Based%20Activation). The main dif-

ference with inetd-style services is that shepherd hands over the listening socket(s) to the daemon; the daemon is then responsible for accepting incoming connections. A handful of environment variables are set in the daemon’s execution environment (see below), which usually checks them using the libsystemd or libelogind client library helper functions (https://

www.freedesktop.org/software/systemd/man/sd_listen_fds.html). The constructor

and destructor for systemd-style daemons are described below.

[Procedure] make-systemd-constructor command endpoints [#:lazy-start? #t]

[#:user #f] [#:group #f] [#:supplementary-groups ’()] [#:directory (default-service-directory)] [#:file-creation-mask #f] [#:create-session? #t] [#:resource-limits ’()] [#:environment-variables (default-environment-variables)]

Return a procedure that starts command, a program and list of argument, as a systemd-style service listening on endpoints, a list of <endpoint> objects. command is started on demand on the first connection attempt on one of endpoints when lazy-start? is true; otherwise it is started as soon as possible. It is passed the listening sockets for endpoints in file descriptors 3 and above; as such, it is equivalent

to an Accept=no systemd socket unit (https://www.freedesktop.org/software/

systemd/man/systemd.socket.html). The following environment variables are set in its environment:

LISTEN_PID

It is set to the PID of the newly spawned process.

LISTEN_FDS

It contains the number of sockets available starting from file descriptor 3—i.e., the length of endpoints.

     LISTEN_FDNAMES

The colon-separated list of endpoint names.

     This must be paired with make-systemd-destructor.
[Procedure] make-systemd-destructor

Return a procedure that terminates a systemd-style service as created by

     make-systemd-constructor.

The following constructor/destructor pair is also available for your convenience, but we

recommend using make-forkexec-constructor and make-kill-destructor instead (this

is typically more robust than going through the shell):

[Procedure] make-system-constructor command. . .

The returned procedure will execute command in a shell and return #t if execution was successful, otherwise #f. For convenience, it takes multiple arguments which will be concatenated first.

[Procedure] make-system-destructor command. . .

Similar to make-system-constructor, but returns #f if execution of the command was successful, #t if not.

4.5 The root Service

The service root is special, because it is used to control the Shepherd itself. It has an alias shepherd. It provides the following actions (in addition to enable, disable and restart which do not make sense here).

status Displays which services are started and which ones are not.

detailed-status

Displays detailed information about every registered service.

load file Evaluate in the shepherd process the Scheme code in file, in a fresh module that uses the (shepherd services) module—as with the --config option of shepherd (see Section 3.1 [Invoking shepherd], page 5).

eval exp Likewise, evaluate Scheme expression exp in a fresh module with all the neces-

sary bindings. Here is a couple of examples:

               # herd eval root "(+ 2 2)"
               4
               # herd eval root "(getpid)"
               1
               # herd eval root "(lookup-running 'xorg-server)"
               (service (version 0) (provides (xorg-server)) ...)
unload service-name

Attempt to remove the service identified by service-name. shepherd will first stop the service, if necessary, and then remove it from the list of registered services. Any services depending upon service-name will be stopped as part of this process.

If service-name simply does not exist, output a warning and do nothing. If it exists, but is provided by several services, output a warning and do nothing. This latter case might occur for instance with the fictional service web-server, which might be provided by both apache and nginx. If service-name is the special string and all, attempt to remove all services except for the Shepherd itself.

reload file-name

Unload all known optional services using unload’s all option, then load filename using load functionality. If file-name does not exist or load encounters an error, you may end up with no defined services. As these can be reloaded at a later stage this is not considered a problem. If the unload stage fails, reload will not attempt to load file-name.

daemonize

Fork and go into the background. This should be called before respawnable services are started, as otherwise we would not get the SIGCHLD signals when they terminate.

4.6 Legacy GOOPS Interface

From its inception in 2002 with negative version numbers (really!) up to version 0.9.x included, the Shepherd’s service interface used GOOPS, Guile’s object-oriented programming system (see Section “GOOPS” in GNU Guile Reference Manual).

There was a <service> class whose instances you could access directly with slot-ref; generic functions such as start and stop had one method accepting a service object and another method accepting the name of a service as a symbol, which it would transparently resolve.

This interface is deprecated. It is still supported in version 0.10.4 but will be removed in the next stable series.

Fortunately, common idioms are easily converted to the current interface. For example, you would previously create a service like so:

(make <service>       ;deprecated GOOPS interface
#:provides '(something)
#:requires '(another thing)
#:start ...
#:stop ...
#:respawn? #t)

With the new interface (see Section 4.1 [Defining Services], page 8), you would write something very similar:

(service '(something)
#:requirement '(another thing)
#:start ...
#:stop ...
#:respawn? #t)

Likewise, instead of writing:

(start 'whatever)

... you would write:

     (start-service (lookup-service 'whatever))

When using one of the deprecated methods, a deprecation warning is emitted, depending on the value of the GUILE_WARN_DEPRECATED environment variable (see Section “Environment Variables” in GNU Guile Reference Manual).

4.7 Service Examples

The configuration file of the shepherd command (see Section 3.1 [Invoking shepherd], page 5) defines, registers, and possibly starts services. Each service specifies other services it depends on and how it is started and stopped. The configuration file contains Scheme code that uses the programming interface of the (shepherd service) module (see Chapter 4 [Services], page 8).

Let’s assume you want to define and register a service to start mcron, the daemon that periodically executes jobs in the background (see Section “Introduction” in GNU mcron Manual). That service is started by invoking the mcron command, after which shepherd should monitor the running process, possibly re-spawning it if it stops unexpectedly. Here’s the configuration file for this one service:

     (define mcron
       (service
         '(mcron)
         ;; Run /usr/bin/mcron without any command-line arguments.
         #:start (make-forkexec-constructor '("/usr/bin/mcron"))
         #:stop (make-kill-destructor)
         #:respawn? #t))
     (register-services (list mcron))

You can write the snippet above in the default configuration file—

~/.config/shepherd/init.scm if you run shepherd as an unprivileged user.

When you launch it, shepherd will evaluate that configuration; thus it will define and register the mcron service, but it will not start it. To start the service, run:

     herd start mcron

Alternatively, if you want mcron to be started automatically when shepherd starts, you can add this snippet at the end of the configuration file:

     (start-in-the-background '(mcron))

Now let’s take another example: sshd, the secure shell daemon of the OpenSSH project (https://www.openssh.com). We will pass sshd the -D option so that it does not “detach”, making it easy for shepherd to monitor its process; we also tell shepherd to check its PID file to determine once it has started and is ready to accept connections:

     (define sshd
       (service
         '(sshd ssh-daemon)  ;for convenience, give it two names
         #:start (make-forkexec-constructor
                  '("/usr/sbin/sshd" "-D")
                  #:pid-file "/etc/ssh/sshd.pid")

         #:stop (make-kill-destructor)
         #:respawn? #t))
     (register-services (list sshd))
     (start-in-the-background '(sshd))

Alternatively, we can start sshd in inetd mode: in that case, shepherd listens for connection and spawns sshd only upon incoming connections. The inetd mode is enabled by passing the -i command-line option:

     (define sshd
       (service
         '(sshd ssh-daemon)
         #:start (make-inetd-constructor
                  '("/usr/sbin/sshd" "-D" "-i")
                  (list (endpoint
                         (make-socket-address AF_INET INADDR_ANY 22))
                        (endpoint
                         (make-socket-address AF_INET6 IN6ADDR_ANY 22)))
                  #:max-connections 10)
         #:stop (make-inetd-destructor)
         #:respawn? #t))
     (register-services (list sshd))
     (start-in-the-background '(sshd))

The make-socket-address procedure calls above return the listening addresses (see Section “Network Socket Address” in GNU Guile Reference Manual). In this case, it specifies that shepherd will accept connections coming from any network interface (“0.0.0.0” in IPv4 notation and “::0” for IPv6) on port 22. The endpoint calls wrap these addresses in endpoint records (see [endpoints], page 16). When a client connects, shepherd accepts it and spawns sshd -D -i as a new transient service, passing it the client connection. The #:max-connections parameter instructs shepherd to accept at most 10 simultaneous client connections.

In these examples, we haven’t discussed dependencies among services—the #:requires keyword of <service>—nor did we discuss systemd-style services. These are extensions of what we’ve seen so far. See Chapter 4 [Services], page 8, for details.

If you use Guix System, you will see that it contains a wealth of Shepherd service definitions. The nice thing is that those give you a complete view of what goes into the service—not just how the service is started and stopped, but also what software package is used and what configuration file is provided. See Section “Shepherd Services” in GNU Guix Reference Manual, for more info.

4.8 Managing User Services

The Shepherd can be used to manage services for an unprivileged user. First, you may want to ensure it is up and running every time you log in. One way to accomplish that is by adding the following lines to ~/.bash_profile (see Section “Bash Startup Files” in The GNU Bash Reference Manual):

if [[ ! -S ${XDG_RUNTIME_DIR-$HOME/.cache}/shepherd/socket ]]; then
shepherd
fi

Then, we suggest the following top-level $XDG_CONFIG_HOME/shepherd/init.scm file, which will automatically load individual service definitions from ~/.config/shepherd/init.d:

(use-modules (shepherd service)
             ((ice-9 ftw) #:select (scandir)))
;; Send shepherd into the background
(perform-service-action root-service 'daemonize)
;; Load all the files in the directory 'init.d' with a suffix '.scm'.
(for-each
  (lambda (file)
    (load (string-append "init.d/" file)))
  (scandir (string-append (dirname (current-filename)) "/init.d")
           (lambda (file)
             (string-suffix? ".scm" file))))

Then, individual user services can be put in $XDG_CONFIG_HOME/shepherd/init.d/,

e.g., for ssh-agent.
;; Add to your ~/.bash_profile:
;;
;; SSH_AUTH_SOCK=${XDG_RUNTIME_DIR-$HOME/.cache}/ssh-agent/socket
;; export SSH_AUTH_SOCK
(use-modules (shepherd support))
(define ssh-agent
  (service
    '(ssh-agent)
    #:documentation "Run `ssh-agent'"
    #:start (lambda ()
              (let ((socket-dir (string-append %user-runtime-dir "/ssh-agent")))
                (unless (file-exists? socket-dir)
                  (mkdir-p socket-dir)
                  (chmod socket-dir #o700))
                (fork+exec-command
                 `("ssh-agent" "-D" "-a" ,(string-append socket-dir "/socket"))
                 #:log-file (string-append %user-log-dir "/ssh-agent.log"))))
    #:stop (make-kill-destructor)
    #:respawn? #t))
(register-services (list ssh-agent))
(start-service ssh-agent)

5 Service Collection

The Shepherd comes with a collection of services that let you control it or otherwise extend its functionality. This chapter documents them.

5.1 Monitoring Service

The monitoring service, as its name suggests, monitors resource usage of the shepherd daemon. It does so by periodically logging information about key resources: heap size (memory usage), open file descriptors, and so on. It is a simple and useful way to check whether resource usage remains under control.

To use it, a simple configuration file that uses this service and nothing else would look like this:

     (use-modules (shepherd service monitoring))
     (register-services
       ;; Create a monitoring service that logs every 15 minutes.
       (list (monitoring-service #:period (* 15 60))))
     ;; Start it!
     (start-service (lookup-service 'monitoring))

Using the herd command, you can get immediate resource usage logging:

     $ herd log monitoring
     service names: 3; heap: 8.77 MiB; file descriptors: 20

You can also change the logging period; for instance, here is how you’d change it to 30 minutes:

     $ herd period monitoring 30

The (shepherd service monitoring) module exports the following bindings:

[Procedure] monitoring-service [#:period (default-monitoring-period)]

Return a service that will monitor shepherd resource usage by printing it every period seconds.

[Variable] default-monitoring-period

This parameter specifies the default monitoring period, in seconds.

5.2 Read-Eval-Print Loop Service

Scheme wouldn’t be Scheme without support for live hacking, and your favorite service manager had to support it too! The REPL service provides a read-eval-print loop (REPL) that lets you interact with it from the comfort of the Guile REPL (see Section “Running Guile Interactively” in GNU Guile Reference Manual).

The service listens for connections on a Unix-domain socket—by default

/var/run/shepherd/repl when running as root and /run/user/uid/shepherd/repl

otherwise—and spawns a new service for each client connection. Clients can use the REPL as they would do with a “normal” REPL, except that it lets them inspect and modify the state of the shepherd process itself.

Caveat: The live REPL is a powerful tool in support of live hacking and debugging, but it’s also a dangerous one: depending on the code you execute, you could lock the shepherd process, make it crash, or who knows what. One particular aspect to keep in mind is that shepherd currently uses Fibers in such a way that scheduling among fibers is cooperative and non-preemptive. Beware!

A configuration file that enables the REPL service looks like this:

(use-modules (shepherd service repl))
(register-services (list (repl-service)))

With that in place, you can later start the REPL:

herd start repl

From there you can connect to the REPL socket. If you use Emacs, you might fancy doing it with Geiser’s geiser-connect-local function (see Geiser User Manual).

The (shepherd service repl) module exports the following bindings.

[Procedure] repl-service [socket-file]

Return a REPL service that listens to socket-file.

[Variable] default-repl-socket-file

This parameter specifies the socket file name repl-service uses by default.

6 Misc Facilities

This is a list of facilities which are available to code running inside of the Shepherd and is considered generally useful, but is not directly related to one of the other topic covered in this manual.

6.1 Errors

[macro] assert expr

If expr yields #f, display an appropriate error message and throw an

     assertion-failed exception.

[Procedure] caught-error key args

Tell the Shepherd that a key error with args has occurred. This is the simplest way to cause caught error result in uniformly formatted warning messages. The current implementation is not very good, though.

[macro] without-system-error expr. . .

Evaluates the exprs, not going further if a system error occurs, but also doing nothing about it.

6.2 Communication

The (shepherd comm) module provides primitives that allow clients such as herd to connect to shepherd and send it commands to control or change its behavior (see Section 4.1 [Defining Services], page 8).

Currently, clients may only send commands, represented by the <shepherd-command> type. Each command specifies a service it applies to, an action name, a list of strings to be used as arguments, and a working directory. Commands are instantiated with

shepherd-command:
[Procedure] shepherd-command action service [#:arguments ’()]

[#:directory (getcwd)]

Return a new command (a <shepherd-command>) object for action on service.

Commands may then be written to or read from a communication channel with the following procedures:

[Procedure] write-command command port

Write command to port.

[Procedure] read-command port

Receive a command from port and return it.

In practice, communication with shepherd takes place over a Unix-domain socket, as discussed earlier (see Section 3.1 [Invoking shepherd], page 5). Clients may open a connection with the procedure below.

[Procedure] open-connection [file]

Open a connection to the daemon, using the Unix-domain socket at file, and return the socket. When file is omitted, the default socket is used.

The daemon writes output to be logged or passed to the currently-connected client using

local-output:

[Procedure] local-output format-string . args

This procedure should be used for all output operations in the Shepherd. It outputs the args according to the format-string, then inserts a newline. It writes to whatever is the main output target of the Shepherd, which might be multiple at the same time in future versions.

Under the hood, write-command and read-command write/read commands as s-expressions (sexps). Each sexp is intelligible and specifies a protocol version. The idea is that users can write their own clients rather than having to invoke herd. For instance, when you type herd status, what is sent over the wire is the following sexp:

(shepherd-command
  (version 0)
  (action status) (service root)
  (arguments ()) (directory "/data/src/dmd"))

The reply is also an sexp, along these lines:

(reply (version 0)
       (result (((service ...) ...)))
       (error #f) (messages ()))

This reply indicates that the status action was successful, because error is #f, and gives a list of sexps denoting the status of services as its result. The messages field is a possibly-empty list of strings meant to be displayed as is to the user.

7 Internals

This chapter contains information about the design and the implementation details of the Shepherd for people who want to hack it.

The GNU Shepherd is developed by a group of people in connection with Guix System (https://www.gnu.org/software/guix/), GNU’s advanced distribution, but it can be used on other distros as well. You’re very much welcome to join us! You can report bugs

to bug-guix@gnu.org and send patches or suggestions to guix-devel@gnu.org.

7.1 Coding Standards

About formatting: Use common sense and GNU Emacs (which actually is the same, of course), and you almost can’t get the formatting wrong. Formatting should be as in Guile and Guix, basically. See Section “Coding Style” in GNU Guix Reference Manual, for more info.

7.2 Design Decisions

Note: This section was written by Wolfgang Jährling back in 2003 and documents the original design of what was then known as GNU dmd. The main ideas remain valid but some implementation details and goals have changed.

The general idea of a service manager that uses dependencies, similar to those of a Makefile, came from the developers of the GNU Hurd, but as few people are satisfied with System V Init, many other people had the same idea independently. Nevertheless, the Shepherd was written with the goal of becoming a replacement for System V Init on GNU/Hurd, which was one of the reasons for choosing the extension language of the GNU project, Guile, for implementation (another reason being that it makes it just so much easier).

The runlevel concept (i.e. thinking in groups of services) is sometimes useful, but often one also wants to operate on single services. System V Init makes this hard: While you can start and stop a service, init will not know about it, and use the runlevel configuration as its source of information, opening the door for inconsistencies (which fortunately are not a practical problem usually). In the Shepherd, this was avoided by having a central entity that is responsible for starting and stopping the services, which therefore knows which services are actually started (if not completely improperly used, but that is a requirement which is impossible to avoid anyway). While runlevels are not implemented yet, it is clear that they will sit on top of the service concept, i.e. runlevels will merely be an optional extension that the service concept does not rely on. This also makes changes in the runlevel design easier when it may become necessary.

The consequence of having a daemon running that controls the services is that we need another program as user interface which communicates with the daemon. Fortunately, this makes the commands necessary for controlling services pretty short and intuitive, and gives the additional bonus of adding some more flexibility. For example, it is easiely possible to grant password-protected control over certain services to unprivileged users, if desired.

An essential aspect of the design of the Shepherd (which was already mentioned above) is that it should always know exactly what is happening, i.e. which services are started and stopped. The alternative would have been to not use a daemon, but to save the state on the file system, again opening the door for inconsistencies of all sorts. Also, we would have to use a separate program for respawning a service (which just starts the services, waits until it terminates and then starts it again). Killing the program that does the respawning (but not the service that is supposed to be respawned) would cause horrible confusion. My understanding of “The Right Thing” is that this conceptionally limited strategy is exactly what we do not want.

The way dependencies work in the Shepherd took a while to mature, as it was not easy to figure out what is appropriate. I decided to not make it too sophisticated by trying to guess what the user might want just to theoretically fulfill the request we are processing. If something goes wrong, it is usually better to tell the user about the problem and let her fix it, taking care to make finding solutions or workarounds for problems (like a misconfigured service) easy. This way, the user is in control of what happens and we can keep the implementation simple. To make a long story short, we don’t try to be too clever, which is usually a good idea in developing software.

If you wonder why I was giving a “misconfigured service” as an example above, consider the following situation, which actually is a wonderful example for what was said in the previous paragraph: Service X depends on symbol S, which is provided by both A and B. A depends on AA, B depends on BB. AA and BB conflict with each other. The configuration of A contains an error, which will prevent it from starting; no service is running, but we want to start X now. In resolving its dependencies, we first try to start A, which will cause AA to be started. After this is done, the attempt of starting A fails, so we go on to B, but its dependency BB will fail to start because it conflicts with the running service AA. So we fail to provide S, thus X cannot be started. There are several possibilities to deal with this:

• When starting A fails, terminate those services which have been started in order to fulfill its dependencies (directly and indirectly). In case AA was running already, we would not want to terminate it. Well, maybe we would, to avoid the conflict with BB. But even if we would find out somehow that we need to terminate AA to eventually start X, is the user aware of this and wants this to happen (assuming AA was running already)? Probably not, she very likely has assumed that starting A succeeds and thus terminating AA is not necessary. Remember, unrelated (running) services might depend in AA. Even if we ignore this issue, this strategy is not only complicated, but also far from being perfect: Let’s assume starting A succeeds, but X also depends on a service Z, which requires BB. In that case, we would need to detect in the first place that we should not even try to start A, but directly satisfy X’s dependency on S with B.

• We could do it like stated above, but stop AA only if we know we won’t need it anymore (for resolving further dependencies), and start it only when it does not conflict with anything that needs to get started. But should we stop it if it conflicts with something that might get started? (We do not always know for sure what we will start, as starting a service might fail and we want to fall back to a service that also provides the particular required symbol in that case.) I think that either decision will be bad in one case or another, even if this solution is already horribly complicated.

• When we are at it, we could just calculate a desired end-position, and try to get there by starting (and stopping!) services, recalculating what needs to be done whenever starting a service fails, also marking that particular service as unstartable, except if it fails to start because a dependency could not be resolved (or maybe even then?). This is even more complicated. Instead of implementing this and thereby producing code that (a) nobody understands, (b) certainly has a lot of bugs, (c) will be unmaintainable and (d) causes users to panic because they won’t understand what will happen, I decided to do the following instead:

• Just report the error, and let the user fix it (in this case, fix the configuration of A) or work around it (in this case, disable A so that we won’t start AA but directly go on to starting B).

I hope you can agree that the latter solution after all is the best one, because we can be sure to not do something that the user does not want us to do. Software should not run amok. This explanation was very long, but I think it was necessary to justify why the Shepherd uses a very primitive algorithm to resolve dependencies, despite the fact that it could theoretically be a bit more clever in certain situations.

One might argue that it is possible to ask the user if the planned actions are ok with her, and if the plan changes ask again, but especially given that services are supposed to usually work, I see few reasons to make the source code of the Shepherd more complicated than necessary. If you volunteer to write and maintain a more clever strategy (and volunteer to explain it to everyone who wants to understand it), you are welcome to do so, of course…

7.3 Service Internals

Under the hood, each service record has an associated fiber, a lightweight execution thread (see Section “Introduction” in Fibers). This fiber encapsulates all the state of its corresponding service: its status (whether it’s running, stopped, etc.), its “running value” (such as the PID of its associated process), the time at which its status changed, and so on. Procedures that access the state of a service, such as service-status, or that modify it, such as start-service (see Section 4.3 [Interacting with Services], page 12), merely send a message to the service’s associated fiber.

This pattern follows the actor model: each of these per-service fibers is an actor. There are several benefits:

• each actor has a linear control flow that is easy to reason about;

• access and modification of the service state are race-free since they are all handled sequentially by its actor;

• the actor’s code is purely functional, which again makes it easier to reason about it.

There are other actors in the code, such as the service registry (see Section 4.2 [Service Registry], page 12). Fibers are used pervasively throughout the code to achieve concurrency.

Note that Fibers is set up such that the shepherd process has only one POSIX thread (this is mandated by POSIX for processes that call fork, with all its warts), and fibers are scheduled in a cooperative fashion. This means that it is possible to block the shepherd process for instance by running a long computation or by waiting on a socket that is not marked as SOCK_NONBLOCK. Be careful!

We think this programming model makes the code base not only more robust, but also very fun to work with—we hope you’ll enjoy it too!

Appendix A GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index

A

actions of services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 actor model, for services . . . . . . . . . . . . . . . . . . . . . . . 29 assertions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

C

canonical name of services. . . . . . . . . . . . . . . . . . . . . . . 8 configuration file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Configuration file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 configuration file, examples. . . . . . . . . . . . . . . . . . . . . 20 constructor of a service . . . . . . . . . . . . . . . . . . . . . . . . 10 constructors, generation of . . . . . . . . . . . . . . . . . . . . . 14

D

daemon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 daemon controller . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 deco, daemon controller . . . . . . . . . . . . . . . . . . . . . . . . . 5 destructor of a service. . . . . . . . . . . . . . . . . . . . . . . . . . 10 destructors, generation of . . . . . . . . . . . . . . . . . . . . . . 14 disabled service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 dmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

E

endpoints, for inetd services . . . . . . . . . . . . . . . . . . . . 16 endpoints, for systemd services . . . . . . . . . . . . . . . . . 16 evaluating code in shepherd . . . . . . . . . . . . . . . . . . . . 18

F

fallback services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

G

generating constructors . . . . . . . . . . . . . . . . . . . . . . . . 14 generating destructors . . . . . . . . . . . . . . . . . . . . . . . . . 14 GOOPS, legacy interface . . . . . . . . . . . . . . . . . . . . . . . 19 graph of services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

H

herd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 6, 7 herding, of daemons. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

I

inetd mode, example . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 inetd-style services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 insecure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 invoking shepherd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

L

log file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 logging. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

O

on-demand, starting services. . . . . . . . . . . . . . . . 16, 17 one-shot services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

P

prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 protocol, between shepherd and its clients. . . . . . 26

R

read-eval-print loop, REPL. . . . . . . . . . . . . . . . . . . . . 23 relative file names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 REPL, read-eval-print loop. . . . . . . . . . . . . . . . . . . . . 23 replacement, or a service . . . . . . . . . . . . . . . . . . . . . . . 13 respawn delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 15 respawn limit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 16 respawning services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 root service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

S

Scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 security . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 service actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 service manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 service registry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Service status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 shepherd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 shepherd Invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 socket activation, starting services . . . . . . . . . . . . . . 17 socket special file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 special services. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Starting a service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 starting a service. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 starting services, via socket activation . . . . . . . . . . 17 Status (of services) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Stopping a service. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 stopping a service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 syslog. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 system errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 systemd-style services . . . . . . . . . . . . . . . . . . . . . . . . . . 17

Concept Index 39

T

termination of a service’s process . . . . . . . . . . . . . . . 10 transient services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

V

virtual services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Procedure and Macro Index

A

actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 assert. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

C

caught-error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

E

endpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 exec-command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

F

for-each-service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 fork+exec-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

L

local-output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 lookup-running . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 lookup-service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

M

make-forkexec-constructor. . . . . . . . . . . . . . . . . . . 14

make-inetd-constructor . . . . . . . . . . . . . . . . . . . . . . 17 make-inetd-destructor . . . . . . . . . . . . . . . . . . . . . . . 17 make-kill-destructor . . . . . . . . . . . . . . . . . . . . . . . . 14

make-system-constructor . . . . . . . . . . . . . . . . . . . . . 18

make-system-destructor . . . . . . . . . . . . . . . . . . . . . . 18

make-systemd-constructor . . . . . . . . . . . . . . . . . . . . 17
make-systemd-destructor . . . . . . . . . . . . . . . . . . . . . 18

monitoring-service . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

O

one-shot-service? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 open-connection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

P

perform-service-action . . . . . . . . . . . . . . . . . . . . . . 13

R

read-command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 register-services . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 repl-service. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 respawn-service? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

S

service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 service-canonical-name . . . . . . . . . . . . . . . . . . . . . . 11 service-documentation . . . . . . . . . . . . . . . . . . . . . . . 11 service-enabled? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 service-provision . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 service-replacement. . . . . . . . . . . . . . . . . . . . . . . . . . 13 service-requirement. . . . . . . . . . . . . . . . . . . . . . . . . . 11 service-respawn-delay . . . . . . . . . . . . . . . . . . . . . . . 11 service-respawn-limit . . . . . . . . . . . . . . . . . . . . . . . 11 service-respawn-times . . . . . . . . . . . . . . . . . . . . . . . 13 service-running-value . . . . . . . . . . . . . . . . . . . . . . . 13 service-running? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

service-startup-failures . . . . . . . . . . . . . . . . . . . . 13

service-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 service-status-changes . . . . . . . . . . . . . . . . . . . . . . 13 service-stopped? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 shepherd-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 start. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

start-in-the-background . . . . . . . . . . . . . . . . . . . . . 13

start-service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 stop-service. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

T

transient-service? . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

U

unregister-services. . . . . . . . . . . . . . . . . . . . . . . . . . 12

W

without-system-error . . . . . . . . . . . . . . . . . . . . . . . . 25 write-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

Variable Index

D

default-bind-attempts . . . . . . . . . . . . . . . . . . . . . . . 16

default-environment-variables . . . . . . . . . . . . . . 15
default-monitoring-period. . . . . . . . . . . . . . . . . . . 23
default-pid-file-timeout . . . . . . . . . . . . . . . . . . . . 15
default-process-
  termination-grace-period. . . . . . . . . . . . . . . . . . 15

                                                default-repl-socket-file . . . . . . . . . . . . . . . . . . . . 24

default-respawn-delay . . . . . . . . . . . . . . . . . . . . . . . 15 default-respawn-limit . . . . . . . . . . . . . . . . . . . . . . . 16

X

XDG_RUNTIME_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Type Index

<service>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 19 <shepherd-command> . . . . . . . . . . . . . . . . . . . . . . . . . . . 25